OpenStudy (anonymous):

Find the number of real solutions of the quadratic equation with: a = -1, b = 1, c = -9. A. no real solutions B. 1 real solution C. 2 real solutions D. 3 real solutions Use the quadratic formula to solve the following quadratic equation: 8x2 + 2x - 3 = 0. A. x = -7/8 or 3/8 B. x = -11/8 or 9/8 C. x = -3/4 or 1/2 D. x = -1/2 or 3/4

zepdrix (zepdrix):

Do you remember the quadratic formula?\[\huge x=\frac{-\color{green}{b}\pm \sqrt{\color{green}{b}^2-4\color{red}{a}\color{royalblue}{c}}}{2\color{red}{a}}\] And they gave us these values for the first question,\[\large \color{red}{a=-1}\]\[\large \color{green}{b=1}\]\[\large \color{royalblue}{c=-9}\] Did you try plugging these values in yet? :D

OpenStudy (raden):

just check it the value of discriminant (D) = b^2 - 4ac. if D > 0 then there are 2 real solutions. let's check it given a = -1 , b=1, and c = 9 so, D = b^2 -4ac D = 1^2 - 4(-1)(9) D = 1 + 36 D = 37 37 > 0 therefore, the quadratic fuction has 2 real solutions